OpenStudy (anonymous):

Solve the system of inequalities by graphing. y x – 2 –4x + y – 3 I really need help with this....

OpenStudy (anonymous):

try putting each inequality in y=mx+b are you familiar with that? so for instance, y= -x-2 is equation 1 y= 4x-3 is equation 2 recall that the slope (rise over run) is the mx variable the b variable tells us where the graph intersects the y-axis Now we graph: Equation 1 start at -2 on the y axis then go down one space, and right one space. place a point. Repeat to extend the line Equation 2 start at -3 on the y axis then rise up 4 spaces, and over 1. place a point. Repeat to extend the line.